OpenStudy (anonymous):

Suppose that y is directly proportional to x and that y=-61 when x=6. Use y=kx to find y when x=6

OpenStudy (anonymous):

y directly proportional to x means : \[y = kx\] Where k = Proportionality constant..

OpenStudy (anonymous):

Now, you have x = 6 and y = -61, so put above and find k from there..

OpenStudy (anonymous):

do I multiply -61*6

OpenStudy (anonymous):

You have : \[y = kx\] Divide by x both the sides to find k here..

OpenStudy (anonymous):

I can divide but how do I do both sides

OpenStudy (anonymous):

You have to find k.. \[\frac{y}{x} = \frac{kx}{x}\] Now cancel whatever you can cancel..

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

On right hand side, can't you cancel x with x?? If yes, then please cancel it..

OpenStudy (anonymous):

I know how to do it but not on the computer on this site

OpenStudy (anonymous):

But you are left with y/x=k

OpenStudy (anonymous):

Then take a notebook and keep doing on that.. Yes you have k = y/x this is right.. Now find k, you have x = 6 and y = -61

OpenStudy (anonymous):

what do I do with x=6 and y=-61

OpenStudy (anonymous):

do I multiply

OpenStudy (anonymous):

k = y/x this means y divided by x I guess..

OpenStudy (anonymous):

so is that the end of it

OpenStudy (anonymous):

10.166

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Did you forget - sign with it??

OpenStudy (anonymous):

yes so it -10.166

OpenStudy (anonymous):

So : \(k = -10.166\) Now, remember this value of k, you will be using this to find y.. Now : \(y = kx\) You have x = 6, here so you will here find y..

OpenStudy (anonymous):

do I divide -10.166 into 6

OpenStudy (anonymous):

No, y = kx means y equals k value multiplied by x..

OpenStudy (anonymous):

Actually, question is I think something for Multiple Choice Question.. Like of 1 mark.. It is simple to guess like question..

OpenStudy (anonymous):

so I multiply -10.166*6

OpenStudy (anonymous):

-60.996

OpenStudy (anonymous):

Or you can say, -61.. See, here you are given with when x = 6, y = -61 And look that in the end question is asking for same that find y when x = 6.. So, basically you have answer within your question..